A researcher gathered a sample of 15 women and 10 men and wanted to know if the amount of women to men was significantly unequal. The researcher computed a Chi-Square and got a result of (report the Chi-Square result and df)
Answer #1

Observed Frequency Expected Frequency

Male 15 25/2 = 12.5

Female 10 25/2 = 12.5

2 (15- 12.5)(10- 12.5)2 12.5 12.5

= 1

So, Chi Square Test Statistic is 1 and df = 2-1 = 1